OpenStudy (anonymous):

A 11 gram sample of a substance that's used to treat thyroid disorders has a k-value of 0.125 Find the substances half-life in days. Round your answer to the nearest tenth.

OpenStudy (welshfella):

5.5 = 11e^(-0.125t) solve for t

OpenStudy (welshfella):

take logs to base e of both sides

OpenStudy (anonymous):

I can't seem to do that on my calculator?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

e^(-0.125t) = 5-5/11 = 0.5 -0.125t = ln 0.5 t = ln 0.5 / -0.125

OpenStudy (anonymous):

5.5? @welshfella

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

@welshfella Thank you!!